OpenStudy (anonymous):

find the x values for -6x^5+10x^4-12x^3+12x^2-10x-6 Please I need help and this is due in 20 mins.

OpenStudy (amistre64):

your question is incomplete

OpenStudy (amistre64):

what does "find the x values" refer to?

OpenStudy (anonymous):

find the x values refers to factoring the equation

OpenStudy (anonymous):

I have no idea how to factor this equation because of how long it is (to the 5th power as the derivative)

OpenStudy (amistre64):

what methods are you aware of?

OpenStudy (amistre64):

and just to be sure, is that last term spose to be -6, or +6 ?

OpenStudy (amistre64):

if it has rational solutions; then we can use the rational roots thrm to gather up a pool of trial and error options.

OpenStudy (anonymous):

-6

OpenStudy (anonymous):

what do you man by what methods?

OpenStudy (amistre64):

methods refer to the ways that people are taught to deal with the question at hand. What ways do you know of that will enable you to factor this?

OpenStudy (amistre64):

this has 1 real root and 4 complex roots so the scope of the solutions would be useful to know as well.

OpenStudy (anonymous):

I know about derivatives, intervals, max/min extrema, quadratic, points of inflection, concave up/concave down

OpenStudy (anonymous):

ahh I guess what I mean is what would be the y intercept and the x intercept/points that can be factored to give a point on the x axis

OpenStudy (amistre64):

the y intercept of course is when x= 0, which leaves us with the last term

OpenStudy (amistre64):

the x intercepts, there is only one of them in terms of real numbers, and that is when y=0 which would make factor it useful

OpenStudy (amistre64):

rational roots test; give all possible combinations of:\[\frac{factors~last\#}{factors~first\#}\]

OpenStudy (amistre64):

if it has any rational roots, itll be from this set: \[\pm\frac{1,2,3,6}{1,2,3,6}=\{\pm\frac{1}{1},\pm\frac{2}{1},\pm\frac{3}{1},\pm\frac{6} {1},\pm\frac{2}{3},\pm\frac{1}{3},\pm\frac{3}{2},\pm\frac{1}{2}\}\]

OpenStudy (anonymous):

so would these be possible x- values then?

OpenStudy (amistre64):

those would be possible x values if this has a root that can be written as a fraction; if it is not one of these, then it would have to be an irrational value

OpenStudy (anonymous):

alright I still dont know how to do this but thanks for your help with evrything I really appreciate it

OpenStudy (amistre64):

do you have a list of options to choose from by chance?

OpenStudy (anonymous):

no the assignment is the analysis and creation of a graph based on the equation given

OpenStudy (amistre64):

then you might wanna run thru the first few derivatives of it; that will at least tell you concavitis and critical values

OpenStudy (amistre64):

i believe the x intercept is an irrational number if you havent made any mistakes in typing the equation in the box

OpenStudy (amistre64):

not decimals, just lazy multiplications \[f = -6x^5+10x^4-12x^3+12x^2-10x-6\] \[f' = -6.5x^4+10.4x^3-12.3x^2+12.2x-10\] \[f'' = -6.5.4x^3+10.4.3x^2-12.3.2x+12.2\]

OpenStudy (anonymous):

yep I found it with my calculator, thanks! Do you by any chance know how to find the x and y intercepts, dumb question but I forgot how to

OpenStudy (amistre64):

y intercept is when x=0; when x=0, all the terms with an x in them become zeros. when adding up a bunch of zeros, they tend to be useless. The y intercept is therefore the term that has no x parts associated with it

OpenStudy (amistre64):

the x intercept for this is not easily found since it is not a number that can be written as a fraction

OpenStudy (amistre64):

we can get closer and closer to the x intercept, and find an approximation for it; but trying to determine an actual value for it is futile http://www.wolframalpha.com/input/?i=-6x%5E5%2B10x%5E4-12x%5E3%2B12x%5E2-10x-6%3B+x+from+-.37+to+-.36